OpenStudy (anonymous):

What are the solutions of 4x2 = −20x?

OpenStudy (ivettef365):

first you move the -20x to the other side by changing the sign to positive so you have 4x^2 + 20 = 0 4x (x+ 5) = 0 4x = 0 and x+ 5 = 0 so what are the 2 values for x

OpenStudy (anonymous):

0 and -5

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

thank u

OpenStudy (ivettef365):

yw :)